# Adaptive Multimodal Person Recognition: A Robust Framework for Handling Missing Modalities

### v2

@@ Abstract
Person identification systems often rely on audio, visual, or behavioral cues, but real-world conditions frequently result in missing or degraded modalities. To address this challenge, we propose a multimodal person identification framework that utilizes gesture as a situational enhancer to supplement traditional modalities like voice and face. Our model employs a unified hybrid fusion strategy, integrating both feature-level and score-level information to maximize representational richness and decision accuracy. Specifically, it leverages multi-task learning to process modalities independently, followed by cross-attention and gated fusion mechanisms. Finally, a confidence-weighted strategy dynamically adapts to missing data, ensuring that our single classification head achieves optimal performance even in unimodal and bimodal scenarios. We evaluate our method on CANDOR, a newly introduced interview-based multimodal dataset, which we benchmark in this work for the first time. Our results demonstrate that the proposed trimodal system achieves 99.51% Top-1 accuracy on person identification tasks. In addition, we evaluate our model on the VoxCeleb1 dataset as a benchmark and reach 99.92% accuracy in bimodal mode, outperforming conventional approaches. Moreover, we show that our system maintains high accuracy even when one or two modalities are unavailable, making it a robust solution for real-world person recognition applications. The code and data for this work are publicly available  [^1] The link has been removed for anonymization reasons..
@@ I Introduction
Typical person identification systems rely heavily on a single modality and risk severe performance degradation whenever that modality fails. This challenge calls for an approach that can effectively leverage the strengths of multiple modalities while downplaying those that are missing or degraded [9, 10]. However, multimodal person identification, which combines cues from several modalities like face, gait, gesture, and voice, can deliver robust and accurate recognition if the modalities are integrated to provide complementary support. Utilizing typical fusion methods [9, 12], which concatenate all features, pose still some challenges in real-world conditions, such as one or more modalities frequently becoming unreliable due to occlusions, poor lighting, background noise, or short utterances [11, 15].
@@
In this work, we utilize gesture information as a situational enhancer to provide supplementary information for person identification, specifically addressing the performance decrements that multimodal systems encounter during modality loss. The proposed system introduces an adaptive fusion strategy that integrates face, gesture, and voice modalities within a unified architecture (more in Section III). The core concepts of our approach to handle modality loss are: confidence weighting, cross-modality bridging and multi-stage fusion with mistake correction.
@@
In confidence weighting each modality is equipped with a confidence network that quantifies the reliability of each modality. Cross-modality bridging is implemented via cross-attention [33], enabling the learnable fusion of a primary modality with multiple auxiliary modalities to produce a more discriminative representation vector. Finally, mistake-correction module rectifies overconfident errors by re-examining logit outputs from the face, gesture, and voice pathways alongside the fused ensemble. The result is a flexible pipeline capable of robust person identification even in substantial modality loss.
@@
To validate our approach, we conducted experiments on the CANDOR dataset [30], providing a new benchmark for this recently introduced interview-style corpus. Furthermore, we used the audiovisual VoxCeleb1 dataset [24] as a benchmark to evaluate our approach compared to previous work. To this end, we utilized the original training, validation, and test splits provided by the dataset. It is essential to note that the multimodal version of VoxCeleb1 consists only of still images, which prevents capturing gesture modalities.
@@
In the following sections, we delve deeper into the specifics of our model and experimental findings. First, we discuss related prior work on modality fusion and highlight how existing methods work with different modalities. Next, we present the proposed trimodal system architecture. We then detail the partitioning of the CANDOR dataset and our data augmentation. Finally, our experimental results and ablation analyses illustrate the robustness of the proposed framework under unimodal, bimodal, and trimodal settings.
@@ II Previous Work
Person identification using multiple modalities has received increasing attention due to the complementary strengths of different input signals. While no prior work has explored the joint use of face, voice, and gesture for person identification, existing studies have investigated various modality combinations, such as face–voice and face-gate fusion. In some cases, gesture-based methods can be related to or overlap with gait-based techniques. In this section, we review related work across these domains, including modality fusion strategies, and summarize reported results on the multimodal VoxCeleb1 dataset as a benchmark.
@@
III Trimodal Person Identification
The proposed framework seeks to characterize each input video using a trimodal representation consisting of face, voice, and gesture. As illustrated in Figure 1, the system initializes three modality-specific pathways using feature vectors extracted from efficient pre-trained encoders. Beginning with the facial pathway, we utilize YOLOv8 [29] to detect and extract faces from the video frames. To ensure optimal input, we employ MagFace [22] for quality assessment, selecting the highest-quality facial frame per video. Finally, an IR101-AdaFace model [8], pre-trained on 12 million images from the WebFace12M dataset is used to extract a 512-dimensional embedding.
For the gesture modality, we utilize TimeSformer [3] to extract 768-dimensional embeddings. We apply the model to 16 equally spaced frames from each video, following the standard input configuration used by TimeSformer for action recognition tasks. Since video lengths range from 4 to 16 seconds, the sampling interval varies between 250 milliseconds and 1 second. To complete the trimodal representation, a ResNet293 [18], pre-trained on the VoxBlink2 [18] and VoxCeleb2 [7] datasets, is used to extract a 256-dimensional voice embedding from each utterance.
III-A Modality-Specific Pathways
@@
where $\mathbf{x}^{(\text{face})}\!\in\mathbb{R}^{512}$ is the output of two dense layers, and $c^{(\text{face})}\!\in(0,1)$ is a scalar confidence value obtained via a small feed-forward head and a Sigmoid $\sigma(\cdot)$ function. This design allows the network to both refine face features and dynamically estimate their reliability, facilitating confidence-weighted fusion in subsequent modules. Moreover, $\mathbf{p}^{(\text{face})}\in\mathbb{R}^{K}$ (with $K$ the number of classes) is logit vector produced by the face-specific classifier head. Analogous expressions hold for gesture- and voice-specific pathways.
III-B Trimodal Cross-Attention
@@
This module bridges the information across the face, gesture, and voice streams, ensuring that each modality interacts with the others rather than being processed in isolation. This interaction is realized through a two-stage process comprising cross-modality projection and cross-attention processing. In the projection stage, each modality receives a projection of the other two modalities into its respective feature space to facilitate alignment.
In the cross-attention stage, each modality-specific pathway acts as the Query ($Q$), while the other two modalities serve as the Key ($K$) and Value ($V$). This configuration allows the target modality to adaptively attend to and aggregate relevant features from the auxiliary streams. For instance, the face cross-attention sub-block utilizes the face embedding as the $Q$ to draw complementary information from the gesture and voice embeddings, which function as the $K$ and $V$.
@@
For instance, the face cross sub-block aggregates gesture and voice embeddings into the face branch:
$\mathbf{z}^{(\text{face})}_{x}$ $=CA_{\text{face}}^{X}\Bigl(\mathbf{z}^{(\text{face})}+\mathrm{proj}_{f}(\mathbf{z}^{(\text{gest})})$
@@
Where $CA_{\text{face}}^{X}(\cdot)$ is the cross-attention block that operates in the face dimension, and $\mathrm{proj}_{f}(\cdot)$ is a learnable linear projection that maps the sum of the gesture and voice embeddings into the face feature space. Analogous cross modules exist for the gesture and voice branches, yielding $\mathbf{z}^{(\text{gest})}_{x}$ and $\mathbf{z}^{(\text{voice})}_{x}$.
III-C Fusion Classification
@@
The fusion gate acts as a dynamic weighting mechanism, selectively emphasizing useful information and filtering out irrelevant noise. It consists of feature concatenation that brings together face, gesture, and voice features into a unified vector $\mathbf{z}_{x}^{\text{concat}}$, then a learnable gating mechanism determines how much each feature dimension contributes.
@@
However, the confidence-weighted fusion mechanism ensures that the contribution of each modality is dynamically adjusted based on confidence scores. Each modality has an associated confidence score ($c^{(\text{face})},c^{(\text{gest})},c^{(\text{voice})}$) which are learned via the confidence network in each modality pathway. The confidence-weighted prediction is computed as:
@@ III-D Decision Making
$\mathbf{p}^{\text{corr}}=\text{MLP}_{\text{corr}}\left(\mathbf{p}^{(\text{face})},\mathbf{p}^{(\text{gest})},\mathbf{p}^{(\text{voice})},\mathbf{p}^{\text{ensemble}}\right).$ (10)
@@
IV Datasets
In this section, we present our procedure for establishing the first benchmark for person identification on the CANDOR dataset. In addition, we use VoxCeleb1 as a standard dataset for multimodal person identification.
IV-A CANDOR Dataset
The CANDOR dataset [30], can serve as a new benchmark automatic speaker recognition. Despite its high potential for multimodal speaker recognition, this dataset has not been widely utilized due to the lack of a standardized partitioning scheme for training and evaluation and the absence of properly segmented data. To address this gap, we have prepared and publicly released the list of segments.
Dataset #Train #Validation #Test #Speakers
CANDOR 87,446 15,632 19,426 1,429
VoxCeleb1 138,361 6,903 8,251 1,251
The CANDOR dataset comprises recordings of 1,450 speakers engaged in discussions on various topics, such as politics, family, and COVID-19, while interacting with different individuals they are meeting for the first time. Audio and video are recorded throughout these conversations, capturing the speakers’ speech, facial expressions, and gestures. As illustrated in Figure 2, speaker participation across sessions varies significantly; while approximately half of the participants appear in only a single session, others are featured in as many as eight. We excluded 21 speakers from our evaluation due to unusable video data caused by disabled cameras, physical obstructions, or the wearing of masks.
Ultimately, the dataset includes 1,429 speakers, each with 16 to 50 training utterances, and 8 to 40 validation and test utterances, stored in MP4 format to extract audio, facial, and gesture features. The number of utterances in each set is reported in the Table I.
IV-B VoxCeleb1 Dataset
Although VoxCeleb1 is categorized as a multimodal dataset, it provides only still facial images extracted from video frames, limiting its capability to capture gesture or temporal visual features.
IV-C Data Augmentation
To enhance the robustness of our speaker recognition model, we employ a multimodal data augmentation strategy. Each modality, which consists of face, gesture, and voice, is augmented independently using a combination of Gaussian noise, dropout regularization, and feature masking.
@@ IV-C Data Augmentation
V Experimental Result
In this study, we developed systems under bimodal and trimodal configurations and conducted evaluations across the full range of unimodal, bimodal, and trimodal scenarios. For the bimodal setup, we utilized voice and face modalities sourced from the VoxCeleb1 and CANDOR datasets. In contrast, the trimodal system was trained solely on the CANDOR dataset and incorporates gesture data in addition to face and voice modalities. Finally, to evaluate the contribution of each in the proposed system, we conducted an ablation study. All the models in this paper were trained in 5 to 8 epochs. We employed a cosine learning rate schedule with warmup and AdamW [19] as the optimizer.
Table II presents the results of the three encoders consisting of ResNet293, IR101-Adaface, and TimesFormer, evaluated for each modality to assess their unimodal performance. Each model is fine-tuned using a classifier head on the VoxCeleb1 and CANDOR datasets separately. The results indicate that on the VoxCeleb1 dataset the models for voice and face reached 99.51% and 99.74%, respectively. However, for the CANDOR dataset the baseline models for voice and face achieved Top-1 accuracies of 97.26% and 97.33%, respectively, significantly outperforming the gesture-based model, which obtained a Top-1 accuracy of 75.4%. This suggests that the gesture embedding vector is more challenging to use for person identification than voice and face features, meaning we cannot rely on it as a standalone modality.
@@ V Experimental Result
Utilizing gesture for identity recognition is relatively uncommon. This modality tends to be highly session-dependent, as model attempts to capture temporal patterns that are sensitive to camera angles and environmental settings. To reveal this dependency, we evaluate the performance of all unimodal models under two conditions: the first where training and testing data for each speaker are drawn from the same session, and the second where they are drawn from different sessions.
TABLE III: Top-1 accuracy (%) in person identification for face, gesture, and voice for speakers with the same train and test sessions and speakers with the different sessions
Condition Face Gesture Voice #Speakers
@@
The detailed results of the session-based evaluation are presented in Table III. For the gesture modality, the model achieves a Top-1 accuracy of 99.93% in same-session scenarios; however, this performance drops significantly to 68.15% when evaluated across different sessions, whereas face and voice modalities maintain relatively stable accuracy. This suggests that session-specific factors, including camera angle, speaking style, and background setup, significantly influence gesture-based identity features. Furthermore, cognitive and contextual variables, such as the conversation topic, the speaker’s mood, and interpersonal dynamics with a partner, potentially affect the results. Consequently, while gesture features are highly reliable within a single session, their generalization across sessions is limited.
@@
V-A Bimodal System
In the bimodal system, we employed the face and voice modalities to design a model that leverages the complementary information from each modality. The primary motivation for implementing the bimodal system in this study was to apply our proposed ideas to the VoxCeleb1 dataset, a widely recognized benchmark in this field. Accordingly, we adopted the same architecture presented in Figure 1, with the only modification being the removal of components related to the gesture modality.

### v3

@@ Abstract
Person identification systems often rely on audio, visual, or behavioral cues, but real-world conditions frequently present with missing or degraded modalities. To address this challenge, we propose a multimodal person identification framework incorporating upper-body motion, face, and voice. Experimental results demonstrate that body motion outperforms traditional modalities such as face and voice in within-session evaluations, while serving as a complementary cue that enhances performance in multi-session scenarios. Our model employs a unified hybrid fusion strategy, fusing both feature-level and score-level information to maximize representational richness and decision accuracy. Specifically, it leverages multi-task learning to process modalities independently, followed by cross-attention and gated fusion mechanisms to exploit both unimodal information and cross-modal interactions. Finally, a confidence-weighted strategy and mistake-correction mechanism dynamically adapt to missing data, ensuring that our single classification head achieves optimal performance even in unimodal and bimodal scenarios. We evaluate our method on CANDOR, a newly introduced interview-based multimodal dataset, which we benchmark in this work for the first time. Our results demonstrate that the proposed trimodal system achieves 99.51% Top-1 accuracy on person identification tasks. In addition, we evaluate our model on the VoxCeleb1 dataset as a widely used evaluation protocol and reach 99.92% accuracy in bimodal mode, outperforming conventional approaches. Moreover, we show that our system maintains high accuracy even when one or two modalities are unavailable, making it a robust solution for real-world person recognition applications. The code and data for this work are publicly available  [^1] The link has been removed for anonymization reasons..
@@ I Introduction
Typical person identification systems rely heavily on a single modality and risk severe performance degradation whenever that modality fails. This challenge calls for an approach that can effectively leverage the strengths of multiple modalities while downplaying those that are missing or degraded [9, 10]. However, multimodal person identification, which combines cues from several modalities like face, body motion, and voice, can deliver robust and accurate recognition if the modalities are integrated to provide complementary support. Utilizing typical fusion methods [9, 12], which concatenate all features, pose still some challenges in real-world conditions, such as one or more modalities frequently becoming unreliable due to occlusions, poor lighting, background noise, or short utterances [11, 15].
@@
In this work, we utilize upper body motion information to provide supplementary information for person identification, specifically addressing the performance decrements that multimodal systems encounter during modality loss. The proposed system introduces an adaptive fusion strategy that integrates face, body motion, and voice modalities within a unified architecture (more in Section IV). The core concepts of our approach to handle modality loss are: confidence weighting, cross-modality bridging and multi-stage fusion with mistake correction.
@@
In confidence weighting each modality is equipped with a confidence network that quantifies the reliability of each modality. Cross-modality bridging is implemented via cross-attention [33], enabling the learnable fusion of a primary modality with multiple auxiliary modalities to produce a more discriminative representation vector. Finally, mistake-correction module rectifies overconfident errors by re-examining logit outputs from the face, motion, and voice pathways alongside the fused ensemble. The result is a flexible pipeline capable of robust person identification even in substantial modality loss.
@@
To train and evaluate our approach, we conducted experiments on the CANDOR dataset [30], providing a new benchmark for this recently introduced interview-style corpus. Furthermore, we used the audiovisual VoxCeleb1 dataset [24] as a benchmark to evaluate our approach compared to previous work. To this end, we utilized the original training, validation, and test splits provided by the dataset. It is essential to note that the multimodal version of VoxCeleb1 consists only of still images, which prevents capturing body motion.
@@
In the following sections, we delve deeper into the specifics of our model and experimental findings. First, we discuss related prior work on modality fusion and highlight how existing methods work with different modalities. We then detail the partitioning of the CANDOR dataset and our data augmentation. Next, we present the proposed trimodal system architecture. Finally, our experimental results and ablation analyses illustrate the robustness of the proposed framework under unimodal, bimodal, and trimodal settings.
@@ II Previous Work
Person identification using multiple modalities has received increasing attention due to the complementary strengths of different input signals. While no prior work has explored the joint use of face, voice, and motion for person identification, existing studies have investigated various modality combinations, such as face–voice and face-gate fusion. In some cases, motion-based methods can be related to or overlap with gait-based techniques. In this section, we review related work across these domains, including modality fusion strategies, and summarize reported results on the multimodal VoxCeleb1 dataset as a benchmark.
@@
III Datasets
In this section, we present our procedure for establishing the first benchmark for person identification on the CANDOR dataset. In addition, we use VoxCeleb1 as a standard dataset for audio-visual person identification.
III-A CANDOR Dataset
The CANDOR dataset [30], can serve as a new benchmark for automatic speaker recognition. Despite its high potential for multimodal speaker recognition, this dataset has not been widely utilized due to the lack of a standardized partitioning scheme for training and evaluation and the absence of properly segmented data. To address this gap, we have prepared and publicly released the list of segments.
Dataset #Train #Validation #Test #Speakers
CANDOR 87,446 15,632 19,426 1,429
VoxCeleb1 138,361 6,903 8,251 1,251
The CANDOR dataset comprises recordings of 1,450 speakers seated in front of a webcam and engaged in discussions on various topics, including politics, family, and COVID-19, while interacting with different individuals whom they are meeting for the first time. Audio and video are recorded throughout these conversations, capturing the speakers’ speech, face, and upper body movements. As illustrated in Figure 1, speaker participation across sessions varies significantly; while approximately half of the participants appear in only a single session, others are featured in as many as eight. We excluded 21 speakers from our evaluation due to unusable video data caused by disabled cameras, physical obstructions, or the wearing of masks.
Ultimately, the dataset includes 1,429 speakers, each with 16 to 50 training utterances, and 8 to 40 validation and test utterances, stored in 30fps MP4 format. The number of utterances in each set is reported in the Table I.
III-B VoxCeleb1 Dataset
Although VoxCeleb1 is categorized as a multimodal dataset, it provides only still facial images extracted from video frames, limiting its capability to capture body motion or spatiotemporal features.
IV Trimodal Person Identification
The proposed framework seeks to characterize each input video using a trimodal representation consisting of face, voice, and body motion. As illustrated in Figure 2, the system initializes three modality-specific pathways using feature vectors extracted from efficient pre-trained encoders. Beginning with the facial pathway, we utilize YOLOv8 [29] to detect and extract faces from the video frames. To ensure optimal input, we employ MagFace [22] for quality assessment, selecting the highest-quality facial frame per video. Finally, an IR101-AdaFace model [8], pre-trained on 12 million images from the WebFace12M dataset is used to extract a 512-dimensional embedding.
For body motion, we utilize TimeSformer [3] to extract 768-dimensional spatiotempora vector. We apply the model to 16 equally spaced frames from each video, following the standard input configuration used by TimeSformer for action recognition tasks. Since video lengths range from 4 to 16 seconds, the sampling interval varies between 250 milliseconds and 1 second. To complete the trimodal representation, a ResNet293 [18], pre-trained on the VoxBlink2 [18] and VoxCeleb2 [7] datasets, is used to extract a 256-dimensional voice embedding from each utterance.
IV-A Modality-Specific Pathways
@@
where $\mathbf{x}^{(\text{face})}\!\in\mathbb{R}^{512}$ is the output of two dense layers, and $c^{(\text{face})}\!\in(0,1)$ is a scalar confidence value obtained via a small feed-forward head and a Sigmoid $\sigma(\cdot)$ function. This design allows the network to both refine face features and dynamically estimate their reliability, facilitating confidence-weighted fusion in subsequent modules. Moreover, $\mathbf{p}^{(\text{face})}\in\mathbb{R}^{K}$ (with $K$ the number of classes) is logit vector produced by the face-specific classifier head. Analogous expressions hold for motion- and voice-specific pathways.
IV-B Trimodal Cross-Attention
@@
This module bridges the information across the face, motion, and voice streams, ensuring that each modality interacts with the others rather than being processed in isolation. This interaction is realized through a two-stage process comprising cross-modality projection and cross-attention processing. In the projection stage, each modality receives a projection of the other two modalities into its respective feature space to facilitate alignment.
In the cross-attention stage, each modality-specific pathway acts as the Query ($Q$), while the other two modalities serve as the Key ($K$) and Value ($V$). This configuration allows the target modality to adaptively attend to and aggregate relevant features from the auxiliary streams. For instance, the face cross-attention sub-block utilizes the face embedding as the $Q$ to draw complementary information from the motion and voice embeddings, which function as the $K$ and $V$.
@@
For instance, the face cross sub-block aggregates motion and voice embeddings into the face branch:
$\mathbf{z}^{(\text{face})}_{x}$ $=CA_{\text{face}}^{X}\Bigl(\mathbf{z}^{(\text{face})}+\mathrm{proj}_{f}(\mathbf{z}^{(\text{motion})})$
@@
Where $CA_{\text{face}}^{X}(\cdot)$ is the cross-attention block that operates in the face dimension, and $\mathrm{proj}_{f}(\cdot)$ is a learnable linear projection that maps the sum of the motion and voice embeddings into the face feature space. Analogous cross modules exist for the motion and voice branches, yielding $\mathbf{z}^{(\text{gest})}_{x}$ and $\mathbf{z}^{(\text{voice})}_{x}$.
IV-C Fusion Classification
@@
The fusion gate acts as a dynamic weighting mechanism, selectively emphasizing useful information and filtering out irrelevant noise. It consists of feature concatenation that brings together face, motion, and voice features into a unified vector $\mathbf{z}_{x}^{\text{concat}}$, then a learnable gating mechanism determines how much each feature dimension contributes.
@@
However, the confidence-weighted fusion mechanism ensures that the contribution of each modality is dynamically adjusted based on confidence scores. Each modality has an associated confidence score ($c^{(\text{face})},c^{(\text{motion})},c^{(\text{voice})}$) which are learned via the confidence network in each modality pathway. The confidence-weighted prediction is computed as:
@@ IV-D Decision Making
$\mathbf{p}^{\text{corr}}=\text{MLP}_{\text{corr}}\left(\mathbf{p}^{(\text{face})},\mathbf{p}^{(\text{motion})},\mathbf{p}^{(\text{voice})},\mathbf{p}^{\text{ensemble}}\right).$ (10)
@@
V Experimental Result
In this study, we developed systems under bimodal and trimodal configurations and conducted evaluations across the full range of unimodal, bimodal, and trimodal scenarios. For the bimodal setup, we utilized voice and face modalities sourced from the VoxCeleb1 and CANDOR datasets. In contrast, the trimodal system was trained solely on the CANDOR dataset and incorporates motion data in addition to face and voice modalities. Finally, to evaluate the contribution of each in the proposed system, we conducted an ablation study. All the models in this paper were trained in 5 to 8 epochs. We employed a cosine learning rate schedule with warmup and AdamW [19] as the optimizer. To improve robustness to modality dropout and noisy data, we applied several data augmentation techniques.
V-A Data Augmentation
To enhance the robustness of our speaker recognition model, we employ a multimodal data augmentation strategy. Each modality, which consists of face, body motion, and voice, is augmented independently using a combination of Gaussian noise, dropout regularization, and feature masking.
@@ V-A Data Augmentation
Table II presents the results of the three encoders consisting of ResNet293, IR101-Adaface, and TimesFormer, evaluated for each modality to assess their unimodal performance. Each model is trained with a frozen encoder, updating only the classifier head on the VoxCeleb1 and CANDOR datasets. The results indicate that on the VoxCeleb1 dataset the models for voice and face reached 99.51% and 99.74%, respectively. On the CANDOR dataset, the baseline models achieved Top-1 accuracies of 97.26% for voice and 97.33% for face, while the motion-based model reached 75.4%.
@@
TABLE III: Top-1 accuracy (%) in person identification for face, motion, and voice for speakers with the same train and test sessions and speakers with the different sessions
Condition Face motion Voice #Speakers
@@
Leveraging body motion for speaker recognition is a relatively rare and innovative approach, offering a complementary modality to traditional face and voice features. Motion captures unique temporal patterns that reflect a speaker’s identity, providing rich behavioral cues often overlooked in conventional methods. To investigate the robustness of this modality, we evaluate all unimodal models under two conditions: (1) when training and testing data for each speaker come from the same session, and (2) when they come from different sessions.
The session-based evaluation results are summarized in Table III. Remarkably, the body motion model achieves a Top-1 accuracy of 99.93% in same-session scenarios, surpassing both voice and face modalities and demonstrating the exceptional discriminative power of motion cues for speaker identification. While performance decreases to 68.15% in cross-session evaluations, face and voice modalities remain more stable, highlighting the unique sensitivity of motion to session-specific factors such as camera angle, speaking style, and environmental conditions. Importantly, these results underscore the potential of motion as a highly informative signal within controlled contexts.
@@
V-B Bimodal System
In the bimodal system, we employed the face and voice modalities to design a model that leverages the complementary information from each modality. The primary motivation for implementing the bimodal system in this study was to apply our proposed ideas to the VoxCeleb1 dataset, a widely recognized benchmark in this field. Accordingly, we adopted the same architecture presented in Figure 2, with the only modification being the removal of components related to the body motion.